Calculate the fraction of oxygen molecules moving with a speed between 700 ms-1 and 710 ms-1 at 300K using the following equation.



f(u) du= dN/N= 4Ï€u2 (M/2Ï€ RT)3/2 e-(Mu^2/2RT) du

1
Expert's answer
2021-03-11T08:41:42-0500

Given Data:

Moving with a speed 700 m/s and 710m/s

Temperature(T)=300k

using the following equation

f(u) du= dN/N= 4Ï€u2 (M/2Ï€ RT)3/2 e-(Mu^2/2RT) du

We get-

f(u)=N(u)du/N

=4Ï€(m/2Ï€kT)^(3/2)u2exp(-mu2/2kT)du

u=710-700=10m/s

Now,

f=4π{(32×1.67×10^-27)/(2π×1.38×10^-23×300)}^3/2×7052×exp{-(32×1.67×10^-27×705^2)/(2×1.38×10^-23×300)}×10

=7.43×10^-3

Hence the fraction of oxygen molecule is 7.43×10^-3